OpenStudy (anonymous):

need help with 3 more problems plz help

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i have 3 more math problems to do then i am done..lol

OpenStudy (anonymous):

i dont want to keep u up

OpenStudy (anonymous):

it ok I'm free

OpenStudy (anonymous):

bonus #1

OpenStudy (anonymous):

i did number 25 already

OpenStudy (anonymous):

ok u want do bonus question 1?

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

3x^2+5x-1=0 x=(-5+-sqrt(25+4))/6 now simplify

OpenStudy (anonymous):

\[\frac{-5\pm \sqrt{5^{2}-4(3\times (-1)}}{(2\times3)} =\frac{-5\pm \sqrt{25+12}}{6}\]

OpenStudy (anonymous):

OK , bye

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

i need help with 2 more problems

OpenStudy (anonymous):

Which questions do you need help on?

OpenStudy (anonymous):

number 18 on this one

OpenStudy (anonymous):

Ok opposite sides of a paralleogram are equal in length so make to equations y2 = 49 so y = 7 (answer) Other equation x + 9 = 4x - 27 subtract x from each side 9 = 3x - 27 Add 27 to each side 18 = 3x divide each side by 3 x = 6 *answer()

OpenStudy (anonymous):

can u show me a break down of the first one plz

OpenStudy (anonymous):

The part about paralellograms?

OpenStudy (anonymous):

One side of the parallelogram is y squared and the other is 49 They are equal in length so if we take the sq root of both of them y = 7

OpenStudy (anonymous):

how u got the answer because i dont get it im confused

OpenStudy (anonymous):

oh we are dividing

OpenStudy (anonymous):

No we are finding the square root of 49 and of y squared because they are equal

OpenStudy (anonymous):

wait are we doing number 18 or bonus number 2

OpenStudy (anonymous):

ohh we have to solve twice

OpenStudy (anonymous):

thats where i got confused

OpenStudy (anonymous):

!8 but I can do bonus 2 for you if you wish

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

I'm on it right now

OpenStudy (anonymous):

thanks a bunch